DYNAMICS IN SOLID MOLECULAR HYDROGEN BELOW 4 K

\underline{DAVID T. ANDERSON}, Department of Chemistry, University of Wyoming, Laramie, WY 82071-3838.

Cryogenic solid molecular hydrogen is a unique condensed phase environment due to a variety of quantum mechanical effects such as large amplitude translational zero point motion of the individual H_{2} molecules, the presence of ortho and para nuclear spin manifolds, and that in the solid phase the individual molecules retain good vibrational and rotational quantum numbers. Equally intriguing is that a variety of non-classical dynamics occur within the crystal at liquid helium temperatures. This talk will highlight some of these dynamics as elucidated from high-resolution FTIR spectroscopy of chemically doped parahydrogen crystals. Some examples of the types of phenomena that have been studied are: ortho-H_{2} spin diffusion and cluster growth, nuclear spin relaxation of ortho-H_{2} to para-H_{2} catalyzed by the presence of a closed shell impurity molecule, and the loss of chemically reactive species introduced into the crystal via tunneling reactions with the surrounding H_{2}.